OpenStudy (anonymous):

if you had 1.68 x 10^28 atoms of zinc,how much would it weight?

OpenStudy (anonymous):

For this questin and the one you have posted next, use the following: - Ratio and proportion, where according to Avogadro's constant, 1 mol = 6.02*10^23 atoms. Hence x mol = 1.68*10^28atoms. Find x. - Next, use the formula n=m/Mr, where n is 'mol', m is required mass and Mr is the mass number of the element or compound according to the periodic table. In this formula, you have everything except for m, so making it as subject of the formula, you get m=n*Mr. Got it?